8.A coin is loaded so that heads is three times as likely as the tails. For 3 independent tosses of the coin find the pmf of the total number of heads realised and the probability of realizing at most 2 heads.

9.Supposse that the random variable x has a p.d.f given by f(x)={cx,0<x<1 o ,elsewhere.Find the values of the constant c hence determine m. so that p(x<or equal to m)=1/2


1
Expert's answer
2021-07-13T05:08:48-0400

8. Let "X=" the number of heads realised: "X\sim Bin(n, p)"

Given "n=3, p=0.75, q=0.25"


"P(X=x)=\dbinom{3}{x}(0.75)^x(0.25)^{3-x}, x=0,1,2,3"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 \\ \hline
 p(x) & 1/64 & 9/64 & 27/64 & 27/64 \\
 
\end{array}"

"P(X\leq2)=1-P(X=3)=1-\dfrac{27}{64}=\dfrac{37}{64}"

"=0.578125"


9.


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{1}cxdx"

"=c\big[\dfrac{x^2}{2}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{c}{2}=1=>c=2"

"\displaystyle\int_{0}^{m}2xdx=\dfrac{1}{2}, 0<m<1"

"[x^2]\begin{matrix}
 m \\
 0
\end{matrix}=\dfrac{1}{2}"

"m^2=\dfrac{1}{2}"

Since "0<m<1," we take "m=\dfrac{\sqrt{2}}{2}."